# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v3

@@ 3.4 Boundary Consistency Regularized Multi-Task Learning (BCRM)
Figure 3: Overview of the BoundMatch architecture processing labeled samples, illustrating multi-task learning (MTL) of boundary detection alongside segmentation. The Boundary-Semantic Fusion (BSF) module integrates learned boundary cues from the boundary head into the segmentation head’s features, promoting sharper segmentation boundaries. Concurrently, the Spatial Gradient Fusion (SGF) module refines the boundary predictions by fusing them with the spatial gradient of the segmentation mask ($\nabla M$). This refinement improves the quality of boundary supervision for the labeled loss and is crucial for generating cleaner boundary pseudo-labels used in the consistency regularization process for unlabeled data.
@@ 3.4.1 Notation and Boundary Detection Head
where the weights are defined by $\beta=\nicefrac{{|z^{+}|}}{{|z|}}$, $\quad 1-\beta=\nicefrac{{|z^{-}|}}{{|z|}}$, with $|z^{+}|$ and $|z^{-}|$ representing the number of boundary and non-boundary pixels, respectively. This reweighting addresses the class imbalance between boundary and non-boundary pixels, ensuring that the model effectively learns to detect boundaries.
@@
While our primary auxiliary task is semantic boundary detection, we also explore the use of binary boundaries in Sec. 4.3.3. We opt for semantic boundaries because they provide richer, class-specific edge information compared to generic binary edges. This detailed supervision can encourage the model to learn features that better distinguish between adjacent classes, potentially leading to improved feature representations in the backbone [11].
@@ 3.4.2 MTL for SS-SS with Boundaries
For the labeled set, the boundary detection head is trained alongside the segmentation head. For unlabeled data, we perform consistency regularization analogous to segmentation as shown in Fig. 1. We generate boundary pseudo-labels from the teacher’s weakly augmented predictions $q^{T,w}$, where $q^{T,w}=q_{\text{fuse}}^{T,w}$ represents the teacher’s fused boundary output (as described in Sec. 3.4.1). These are converted to hard pseudo-labels using threshold $\tau_{bdry}$: $\hat{q}^{T}=\mathbf{1}\Big{(}q^{T,w}\geq\tau_{bdry}\Big{)}$. The unsupervised boundary loss then enforces consistency between these pseudo-labels and the student’s strongly augmented predictions:
@@
Note that when SGF is enabled (Sec. 3.5.2), we use the refined boundary prediction $q_{\text{refine}}^{T,w}$ instead of $q_{\text{fuse}}^{T,w}$ to generate higher-quality pseudo-labels, as the refinement process produces cleaner boundaries by incorporating spatial gradient information.
@@ 3.5 Boundary-Aware Modeling and Refinement
To further enhance the boundary quality and boost segmentation performance, we propose two complementary modules: Boundary-Semantic Fusion (BSF) and Spatial Gradient Fusion (SGF). BSF and SGF are simple yet effective mechanisms that refine the boundary and segmentation predictions by incorporating boundary cues and spatial gradients as shown in Fig. 3.
@@ 3.5.1 Boundary-Semantic Fusion (BSF)
Inspired by GSCNN [10], BSF explicitly integrates boundary predictions into the segmentation head. The output from the boundary detection head is concatenated with the features from the Atrous Spatial Pyramid Pooling (ASPP) using the bottleneck layer of the DeepLabV3+ decoder. We emphasize that there are no extra layers introduced by BSF, and we only add extra channels for the bottleneck layer. This explicit incorporation of boundary cues helps localize object boundaries more accurately in the final segmentation.
@@ 3.5.2 Spatial Gradient Fusion (SGF)
Semantic boundary detection may suffer from coarse predictions. To refine the boundaries, we adopt the spatial gradient operator as in RPCNet [27]. The spatial gradient of the semantic mask $M$ is computed as:
@@
where $\operatorname{pool}$ denotes adaptive average pooling with a $3\times 3$ kernel. An example of the spatial gradient operator is shown in Fig. 4. The resulting gradient map, $q_{\nabla M}$, is concatenated with the boundary prediction $q_{\text{fuse}}$ for each channel:
$[q_{\text{fuse},0},\,q_{\nabla M,0},\,q_{\text{fuse},1},\,q_{\nabla M,1},\,\dots,\,q_{\text{fuse},C-1},\,q_{\nabla M,C-1}],$ (9)
forming a $2C$-channel activation map. This map is then processed by a convolutional layer with $C$-grouped convolutions to yield a refined semantic boundary probability map $q_{\text{refine}}$. This grouped convolution refines the boundary predictions by considering the spatial gradients, effectively sharpening the boundaries between different semantic regions. The refined boundaries are supervised using the same $H_{bdry}$ loss with the labeled set.
@@ 3.6 BoundMatch
Although demonstrated primarily with our SAMTH baseline throughout the experiments, BoundMatch is designed to be flexible and compatible with various CR methods. For instance, BoundMatch can be integrated into UniMatch [9], a single-model CR method, by incorporating the boundary consistency loss alongside UniMatch’s existing feature-level and strong-view consistency losses. We observed that using only a single model (without a teacher) tends to generate noisy boundary pseudo-labels; therefore, employing an EMA-updated teacher model is crucial for stability and quality.
Similarly, BoundMatch can also be integrated with PrevMatch [44], a state-of-the-art (SOTA) method based on UniMatch, by computing the boundary consistency loss using its robust ensemble pseudo-labels, thereby circumventing potential instability associated with single-iteration predictions.
Furthermore, the architectural simplicity of the BSF and SGF fusion modules facilitates the incorporation of BoundMatch’s boundary-enhancement components into diverse segmentation models, a versatility explored in our ablation studies.
@@ 4.1.1 Datasets
PASCAL VOC 2012 [55] is a standard object recognition benchmark consisting of 10,582 images annotated for 21 classes (20 object categories plus background). Following established SS-SS protocols, we consider two training settings: the “Classic” protocol uses only the high-quality subset of 1,464 images for the labeled pool, while the “Blender” protocol samples labeled images randomly from the entire dataset.
@@ 4.1.2 Implementation Details
Our experiments utilize ResNet-50 and ResNet-101 as encoder backbones across all datasets, coupled with a DeepLabV3+ segmentation head using an output stride of 16 for training efficiency. For generating semantic and binary boundary labels, we employ the on-the-fly (OTF) strategy from [11]. This approach integrates boundary extraction into the data loading pipeline, guaranteeing consistent boundary widths despite random resizing augmentations.
@@
4.2 Comparisons with State-of-the-Art Methods
4.2.1 Cityscapes
Tab. 1 summarizes the quantitative results on the Cityscapes dataset across various labeled data splits ($\nicefrac{{1}}{{16}}$, $\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{4}}$) for both ResNet-50 and ResNet-101 backbones. Compared with recent state-of-the-art methods, SAMTH+BoundMatch achieves significant performance improvements in the low-data regime. Notably, incorporating the boundary-aware components allows SAMTH+BoundMatch to consistently outperform the SAMTH baseline and achieves 1.4% and 2.4% improvements on the difficult $\nicefrac{{1}}{{16}}$ split with ResNet-50 and ResNet-101 respectively. We attribute this improvement to the additional regularization provided by the boundary multi-task learning objective, as strong regularization techniques are often particularly effective in low-supervision scenarios. The qualitative results presented in Fig. 5 further illustrate that SAMTH+BoundMatch produces more precise object boundaries and improves segmentation accuracy for challenging classes such as “wall”, “pole”, “train”, and “bus”. Our proposed simple baseline, SAMTH, also demonstrates strong performance, achieving competitive results against previous SOTA methods. Overall, these findings establish SAMTH as a strong baseline for urban scene SS-SS, while SAMTH+BoundMatch demonstrates competitive performance against state-of-the-art methods through its effective integration of boundary awareness.
As stated previously, BoundMatch can be integrated with other SS-SS methods, and we have shown results for UniMatch [9] and PrevMatch [44] in Tab. 1. For a fair comparison, we reproduced UniMatch and PrevMatch under the same experimental settings as our SAMTH baseline. Incorporating BoundMatch into these methods, we observe consistent performance gains across all splits and backbones. We believe that the performance gains are not as significant as the gains achieved by SAMTH+BoundMatch because UniMatch and PrevMatch already leverage strong consistency regularization techniques which incurs multiple loss functions and may reduce the benefits of the boundary multi-task learning. In future works, we will explore approaches to better balance the regularization signals, but we believe that the results shown here proves the effectiveness of BoundMatch in enhancing SS-SS methods.
[…]
@@ 4.2.1 Cityscapes
CVCC [32] [CVPR’23] $74.9$ $76.4$ $77.3$
iMAS [38] [CVPR’23] $74.3$ $77.4$ $78.1$
@@
UCCL [36] [ICASSP’25] $75.8$ $77.2$ $78.2$
NRCR [45] [NN’25] - $77.0$ $77.9$ $78.2$ $79.5$
@@
4.2.2 BDD100K, SYNTHIA, and ADE20K
Table 2: Comparison of SAMTH + BoundMatch with UniMatch on three datasets using DeepLabV3$+$ (ResNet-50).
@@ 4.2.2 BDD100K, SYNTHIA, and ADE20K
To assess the generalizability of our approach, we extended the evaluation to additional datasets: the urban driving scene datasets BDD100K and SYNTHIA, as well as the challenging academic benchmark ADE20K, known for its diversity and difficulty in semantic segmentation. As shown in Tab. 2, SAMTH+BoundMatch consistently yields significant improvements over supervised baselines and UniMatch on these datasets. For instance, on BDD100K, our method improves the mean IoU by over 3% for the $\nicefrac{{1}}{{64}}$ split. These results demonstrate that our boundary-aware learning framework effectively leverages unlabeled data across various domains, including challenging datasets with complex scenes. For qualitative visualizations, please refer to Fig. 13 in the supplementary material, which illustrates the improved segmentation boundaries and overall performance of SAMTH+BoundMatch compared to UniMatch.
4.2.3 Pascal VOC 2012
We further evaluate our method on the Pascal VOC 2012 dataset using both the Classic and Blender splits, acknowledging the challenges posed by this benchmark. It is important to note that Pascal VOC 2012 boundary annotations are known to be noisy, and boundary regions are typically treated as “ignore labels” during training and evaluation (see Sec. E.2). Therefore, in this benchmark, we hypothesized that the gains from BoundMatch will be limited compared to the previous high-quality datasets. However, experiment results demonstrate that BoundMatch can still provide significant improvements in segmentation performance, especially in the low labeled data regime as shown in this section.
Table 3: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Classic splits. All methods are trained using DeepLabV3$+$ (ResNet-50/101).
@@ 4.2.3 Pascal VOC 2012
Table 4: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Blender splits. All methods are trained using DeepLabV3$+$ (ResNet-50).

### v4

@@ 3.4 Boundary Consistency Regularized Multi-Task Learning (BCRM)
Figure 3: Overview of the BoundMatch architecture processing labeled samples, illustrating multi-task learning (MTL) of boundary detection alongside segmentation. The Boundary-Semantic Fusion (BSF) module integrates learned boundary cues from the boundary head into the segmentation head’s features, promoting sharper segmentation boundaries. Concurrently, the Spatial Gradient Fusion (SGF) module refines the boundary predictions by fusing them with the spatial gradient of the segmentation mask ($\nabla M$). This refinement improves the quality of boundary supervision for the labeled loss and is important for generating cleaner boundary pseudo-labels used in the consistency regularization process for unlabeled data.
@@ 3.4.1 Notation and Boundary Detection Head
where the weights are defined by $\beta=\nicefrac{{|z^{+}|}}{{|z|}}$, $\quad 1-\beta=\nicefrac{{|z^{-}|}}{{|z|}}$, with $|z^{+}|$ and $|z^{-}|$ representing the number of boundary and non-boundary pixels, respectively. This reweighting addresses the class imbalance between boundary and non-boundary pixels, ensuring that the model learns to detect boundaries.
@@
While our primary auxiliary task is semantic boundary detection, we also explore the use of binary boundaries in Sec.˜4.3.3. We opt for semantic boundaries because they provide richer, class-specific edge information compared to generic binary edges. This detailed supervision can encourage the model to learn features that better distinguish between adjacent classes, potentially leading to improved feature representations in the backbone [11].
@@ 3.4.2 MTL for SS-SS with Boundaries
For the labeled set, the boundary detection head is trained alongside the segmentation head. For unlabeled data, we perform consistency regularization analogous to segmentation as shown in Fig.˜1. We generate boundary pseudo-labels from the teacher’s weakly augmented predictions $q^{T,w}$, where $q^{T,w}=q_{\text{fuse}}^{T,w}$ represents the teacher’s fused boundary output (as described in Sec.˜3.4.1). These are converted to hard pseudo-labels using threshold $\tau_{bdry}$: $\hat{q}^{T}=\mathbf{1}\Big(q^{T,w}\geq\tau_{bdry}\Big)$. The unsupervised boundary loss then enforces consistency between these pseudo-labels and the student’s strongly augmented predictions:
@@
Note that when SGF is enabled (Sec.˜3.5.2), we use the refined boundary prediction $q_{\text{refine}}^{T,w}$ instead of $q_{\text{fuse}}^{T,w}$ to generate higher-quality pseudo-labels, as the refinement process produces cleaner boundaries by incorporating spatial gradient information.
From a noise-robustness perspective, BCRM leverages multi-view learning principles where the segmentation and boundary detection heads serve as two complementary views of the same visual understanding task. Unlike methods that derive boundaries directly from segmentation outputs (i.e. BoundaryMatch [15] creates dependent pseudo-labels), our boundary head learns from hierarchical backbone features independently, providing a distinct supervisory signal. This independence is important: when one task branch generates erroneous pseudo-labels, the other branch can continue learning from its own independently-generated pseudo-labels, preventing error propagation between tasks. Following noise-robust learning principles identified in [37], such multi-view consistency with independent learning objectives aims to helps prevent overfitting to task-specific noise patterns.
@@ 3.5 Boundary-Aware Modeling and Refinement
To further enhance the boundary quality and boost segmentation performance, we propose two complementary modules: Boundary-Semantic Fusion (BSF) and Spatial Gradient Fusion (SGF). BSF integrates learned boundary cues into the segmentation decoder to improve object delineation, while SGF refines boundary predictions using spatial gradients from the segmentation mask to produce cleaner boundary pseudo-labels. These simple fusion mechanisms establish bidirectional information flow between tasks as shown in Fig.˜3.
@@ 3.5.1 Boundary-Semantic Fusion (BSF)
Inspired by GSCNN [10], BSF explicitly integrates boundary predictions into the segmentation head through concatenation-based fusion. Formally, let $F_{\text{ASPP}}\in\mathbf{R}^{(C_{\text{ASPP}}\times H/16\times W/16)}$ denote the ASPP features and $q\in\mathbf{R}^{(C\times H/16\times W/16)}$ denote the downsampled boundary features from the boundary prediction head which is detached from gradient computation. The BSF operation is defined as:
$F_{\text{fused}}=\text{Conv}_{1\times 1}([F_{\text{ASPP}};q])$ (8)
where $\cdot$$\cdot$] denotes concatenation along the channel dimension, and $\text{Conv}_{1\times 1}$ is the existing bottleneck layer in the DeepLabV3+ decoder. This design choice preserves full information from both modalities without information loss through element-wise operations. We emphasize that BSF introduces no additional layers; we only expand the input channels of the existing bottleneck layer from $C_{\text{ASPP}}$ to $C_{\text{ASPP}}+C$, maintaining architectural simplicity while enabling effective feature fusion. By incorporating boundary cues, BSF aims to provide the segmentation decoder with complementary edge information that may assist in object delineation.
@@ 3.5.2 Spatial Gradient Fusion (SGF)
While the boundary detection head learns from hierarchical features, it can benefit from geometric cues present in the segmentation predictions. We adopt the spatial gradient operator from RPCNet [27] to extract these cues. The spatial gradient of the semantic mask $M$ is computed as:
@@
where $\operatorname{pool}$ denotes adaptive average pooling with a $3\times 3$ kernel. An example of the spatial gradient operator is shown in Fig.˜4. The resulting gradient map, $q_{\nabla M}$, is concatenated with the boundary prediction $q_{\text{fuse}}$ for each channel:
$q_{\text{fuse},0},\,q_{\nabla M,0},\,q_{\text{fuse},1},\,q_{\nabla M,1},\,\dots,\,q_{\text{fuse},C-1},\,q_{\nabla M,C-1}],$ (10)
forming a $2C$-channel activation map. This interleaving arrangement ensures that each paired boundary-gradient features for the same class are processed together. The map is then processed by a single $C$-grouped convolutional layer:
$q_{\text{refine}}=\text{GroupConv}_{C}([q_{\text{fuse,0}},q_{\nabla M,0},\dots])$ (11)
where $\text{GroupConv}_{C}$ denotes a grouped convolution with $C$-groups. This design choice follows established semantic boundary detection methods [26, 53] where per-class grouped convolutions enable class-specific boundary refinement. Each group independently refines its corresponding class boundary using the paired spatial gradient information, preserving class-specific edge patterns while maintaining computational efficiency through a single convolution operation. The refined boundaries are supervised using the same $H_{bdry}$ loss with the labeled set.
@@ 3.6 BoundMatch
The design of BoundMatch emphasizes modularity and flexibility. The framework’s core principles—boundary CR through independent task heads, bidirectional fusion between tasks, and hierarchical feature extraction—are architecture-agnostic. This modularity enables practitioners to adopt different configurations based on their requirements: full BoundMatch for maximum accuracy, BCRM+SGF for maintaining inference speed, or BCRM alone for minimal overhead. While demonstrated primarily with SAMTH, BoundMatch integrates naturally with other CR methods. For UniMatch [9], we incorporate boundary consistency alongside its existing feature-level and strong-view losses, using an EMA teacher for stable boundary pseudo-labels. For PrevMatch [45], we leverage its robust ensemble pseudo-labels for boundary consistency, circumventing single-iteration instabilities.
BoundMatch’s approach differs from prior methods by learning boundaries independently through dedicated task heads rather than deriving them from segmentation outputs. The teacher model generates boundary pseudo-labels from its learned boundary head (potentially refined by SGF), providing supervision that is complementary to, rather than dependent on, segmentation predictions. This independence aims to provide noise robustness: when segmentation pseudo-labels are uncertain at object edges, the boundary task can provide additional supervision signals, and vice versa. The implementation of BSF and SGF relies on standard operations—feature concatenation and grouped convolutions—which facilitates integration with various segmentation architectures without requiring substantial modifications.
@@ 4.1.1 Datasets
PASCAL VOC 2012 [56] is a standard object recognition benchmark consisting of 10,582 images annotated for 21 classes (20 object categories plus background). Following established SS-SS protocols, we considered two training settings: the “Classic” protocol uses only the high-quality subset of 1,464 images for the labeled pool, while the “Blender” protocol samples labeled images randomly from the entire dataset.
@@ 4.1.2 Implementation Details
Our experiments utilize ResNet-50 and ResNet-101 as encoder backbones across all datasets, coupled with a DeepLabV3+ segmentation head using an output stride of 16 for training efficiency. For generating semantic and binary boundary labels, we employ the on-the-fly (OTF) strategy from [11] by applying signed distance function and using distance of 2 to obtain per-class boundaries. This approach integrates boundary extraction into the data loading pipeline, guaranteeing consistent boundary widths despite random resizing augmentations.
@@
Mean intersection over union (mIoU) is the primary evaluation metric for semantic segmentation, but is known to be insensitive to boundary-specific errors [58, 59]. To quantitatively assess boundary improvements, we adopt Boundary IoU (BIoU) and Boundary F1 Score (BF1), which explicitly measure boundary alignment and detection quality, respectively (see Appendix C.1 for detailed definitions).
@@
We use the same experimental setup for all our ablation studies, in Sec.˜4.3 unless otherwise specified.
Table 2: Comparison with recent state-of-the-art methods on Cityscapes. All methods use DeepLabV3++ with ResNet-50/101 backbones. $\dagger$} denotes method reproduced by us using the same training partition protocol. Bold and underline denotes the best and the second-best results, respectively. Results are averaged over three runs following [9].
@@
CVCC [32] [CVPR’23] $74.9$ $76.4$ $77.3$ - - -
iMAS [39] [CVPR’23] $74.3$ $77.4$ $78.1$ - - -
@@
UCCL [36] [ICASSP’25] $75.8$ $77.2$ $78.2$ - - -
NRCR [37] [NN’25] - $77.0$ $77.9$ - $78.2$ $79.5$
@@
Table 3: Comparison of SAMTH + BoundMatch with UniMatch on three datasets using DeepLabV3++ (ResNet-50).
@@
Table 4: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Classic splits. All methods are trained using DeepLabV3++ (ResNet-50/101).
@@
Table 5: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Blender splits. All methods are trained using DeepLabV3++ (ResNet-50).
@@
4.2 Comparisons with State-of-the-Art Methods
4.2.1 Cityscapes
Tab.˜2 summarizes the quantitative results on the Cityscapes dataset across various labeled data splits ($\nicefrac{{1}}{{16}}$, $\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{4}}$) for both ResNet-50 and ResNet-101 backbones. We have also included boundary metrics in Tab.˜15 in the Appendix. Compared with recent state-of-the-art methods, SAMTH+BoundMatch achieves competitive performances, outperforming most evaluation protocols. Notably, incorporating the boundary-aware components allows SAMTH+BoundMatch to consistently outperform the SAMTH baseline and achieves 1.4% and 2.4% improvements on the difficult $\nicefrac{{1}}{{16}}$ split with ResNet-50 and ResNet-101 respectively. We attribute this improvement to the additional regularization provided by the boundary multi-task learning objective, as strong regularization techniques are often particularly effective in low-supervision scenarios. The qualitative results presented in Fig.˜5 further illustrate that SAMTH+BoundMatch produces more precise object boundaries and improves segmentation accuracy for challenging classes such as “wall”, “pole”, “train”, and “bus”. Our proposed simple baseline, SAMTH, also demonstrates strong performance, achieving competitive results against previous SOTA methods. Overall, these findings establish SAMTH as a strong baseline for urban scene SS-SS, while SAMTH+BoundMatch demonstrates competitive performance against state-of-the-art methods through its effective integration of boundary awareness.
As stated previously, BoundMatch can be integrated with other SS-SS methods, and we have shown results for UniMatch [9] and PrevMatch [45] in Tab.˜2. For a fair comparison, we reproduced UniMatch and PrevMatch under the same experimental settings as our SAMTH baseline. Incorporating BoundMatch into these methods, we observe consistent performance gains across all splits and backbones. However, UniMatch and PrevMatch already leverage strong consistency regularization techniques which incurs multiple loss functions and may reduce the benefits of the boundary multi-task learning. In future works, we will explore approaches to better balance the regularization signals, but we believe that the results shown here proves the effectiveness of BoundMatch in enhancing SS-SS methods.
4.2.2 BDD100K, SYNTHIA, and ADE20K
To assess the generalizability of our approach, we extended the evaluation to additional datasets: the urban driving scene datasets BDD100K and SYNTHIA, as well as the challenging academic benchmark ADE20K, known for its diversity and difficulty in semantic segmentation. As shown in Tabs.˜3 and 18, SAMTH+BoundMatch consistently improves over supervised baselines and UniMatch on these datasets. For instance, on BDD100K, our method improves the mean IoU by over 3% for the $\nicefrac{{1}}{{64}}$ split. These results demonstrate that our boundary-aware learning framework leverages unlabeled data across various domains, including challenging datasets with complex scenes. For qualitative visualizations, please refer to Fig.˜15 in the supplementary material, which illustrates the improved segmentation boundaries and overall performance of SAMTH+BoundMatch compared to UniMatch.
4.2.3 Pascal VOC 2012
We further evaluate our method on the Pascal VOC 2012 dataset using both the Classic and Blender splits, acknowledging the challenges posed by this benchmark. It is important to note that Pascal VOC 2012 boundary annotations are known to be noisy, and boundary regions are typically treated as “ignore labels” during training and evaluation (see Appendix E.3). Therefore, in this benchmark, we hypothesized that the gains from BoundMatch will be limited compared to the previous high-quality datasets. However, experiment results demonstrate that BoundMatch can still provide improvements in segmentation performance, especially in the low labeled data regime as shown in this section. We believe this is due to the strong regularization of using multi-task CR objectives with label-refinement strategy potentially improving noise robustness as theorized in [37].